OpenStudy (anonymous):

(x^7+x^6)(x^7+x^6-x^5) multiply

OpenStudy (anonymous):

x^6* ( x + 1 ) * x^5 * ( x² + x + 1) = x^11 * ( x + 1) * ( x ² + x + 1 ) = x^11 * (x³ + x² + x + x² + x + 1) = x^11 * ( x³ + 2x² + 2x + 1 ) = x^14 + 2x^13 + 2x^12 + x^11

OpenStudy (earthcitizen):

just to confirm, x^14+2 x^13-x^11